Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness Technical integration and coordination
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 617     Entry time: 02 Jun 2025, 09:51
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 02.06.2025 
General Announcement:
- The ongoing delivery sees a HXRSS at 12GeV with 1mJ total intensity (~350uJ background). 

CTRL:
- Nothing major to report.

DA:
- Nothing to be reported.

EEE:
- Nothing to report.

VAC:
- Not present.

XPD:
- No issues to report.
- Big brother stopped SA1 because of HIREX power limit: J. Gruenert reports that all pulses of the train are now being seen by Big Brother and things worked as expected.

DET:
- Nothing major to report.

ITDM: 
- Nothing to report.

XRO:
- Nothing to report.

LAS:
- Ticket was discussed for SPB (timing issues): under investigation and a bug on the code has been found and is under work.

UNSYS:
- Issue with PS in cell#8 in SA1 was already fixed remotely.

SPB:
- Two issues: 1. HIREX watchdog needs a change of filter then works fine; 2. Issue with sample delivery leads that SPB will not use beam tonight (FXE can take if wanted).
- uTCA tested a monitor for received timing info together with ITDM and this test was performed parasitically at SPB.

FXE:
- Nothing to report.

MID:
- Nothing to report.

HED:
- HXRSS performance this week is excellent.

SCS:
- Not present.

SQS:
- Nothing to report.

SXP:
- Not present.
ELOG V3.1.4-7c3fd00